Use the information below for Harding Company to answer the question that follow.

Harding Company
Accounts payable $29,326
Accounts receivable 68,035
Accrued liabilities 6,348
Cash 18,912
Intangible assets 40,462
Inventory 87,761
Long-term investments 90,008
Long-term liabilities 75,920
Marketable securities 36,002
Notes payable (short-term) 22,258
Property, plant, and equipment 624,553
Prepaid expenses 2,032

Based on the data for Harding Company, what is the amount of working capital?
Answer #1

Calculation of Working capital:

Working capital = Current asset - current liability

Current asset = 68,035+18,912+87,761+36,002+2,032 = 212,742

Current liabilities = 29,326+6,348+22,258 = 57,932

Working Capital = 212,742-57,932 = 154,810